OpenStudy (anonymous):

Find the vertex, focus, directrix, and focal width of the parabola. -1/40x^2=y

OpenStudy (anonymous):

answers: Vertex: (0, 0); Focus: (0, -10); Directrix: y = 10; Focal width: 160 Vertex: (0, 0); Focus: (-20, 0); Directrix: x = 10; Focal width: 160 Vertex: (0, 0); Focus: (0, -10); Directrix: y = 10; Focal width: 40 Vertex: (0, 0); Focus: (0, 10); Directrix: y = -10; Focal width: 10

OpenStudy (campbell_st):

well write in the form \[x^2 = 4ay\] so your question becomes \[x^2 = -40y\] so the vertex is at the origin and the parabola is concave down. \[x^2 = - 4 \times 10y\] the value of a, the focal length is 10 and will be above the focus.. so the directrix is y = 10 focus is 10 below the vertex and on the line of symmetry x = 0 (0, -10) so all you need to do is find the width of the parabola...

OpenStudy (anonymous):

so the focal width is 40?

OpenStudy (anonymous):

@campbell_st